your company purchases a machine for $14,000 with a 6-year tax life. the soyd method is used for depreciation, and the tax salvage is zero. Neglect taxes in this part. After the third year of life, the company is thinking about replacing the machine with a new one. it can be sold now for $10,000. Next year it will be worth only $6,000 and in two years, only $4,000. Three years from now the machine will have no resale value. the operating cost of the machine is expected to be constant for the next three years at $1,000 per annum. The new machine has a life of 10 years with a NAC of $5,000. Should the old machine be replaced wih the new one if the company's MARR is 10%? explain
Answer #1

The NAC for 1 year is (10,000 ­- 6000)(A/P, 0.1, 1) + (6000*0.1 )+ 1000

= 4,000 × 1.1 + 600 + 1,000 = $6,000

The NAC for 2 years is (10,000- ­ 4000)(A/P, 0.1, 2) + 4000*0.1 + 1000

  = 6,000 × 0.5762 + 400 + 1,000

= 3,457.20 + 400 + 1,000 = $4,857.20

The NAC for 3 years is (10,000)(A/P, 0.1, 3)

= 4,021 + 1,000

= $5,021.

The economic life of the old machine is two years. It's NAC at this life is less than that of the challenger. Keep the old machine.